### v1

@@ Abstract
Index Terms:
Real-Time Flow Estimation, Embedded FPGAs, Soft Sensor, On-device Inference
@@ I Introduction and Related Work
Combined sewer systems (CSS) are widely used worldwide but face increasing challenges from extreme weather exacerbated by pollution and climate change. These conditions often lead to urban overflow events, significantly impacting public health and environmental sustainability [1]. Real-time control systems offer a promising alternative by preventing CSS overflow through dynamic flow monitoring. However, these systems are hindered by high operational costs and challenges in achieving continuous, accurate monitoring and real-time data processing [2]. Current flow monitoring techniques mainly rely on hydrological formulas, which, while cost-effective, lack the precision necessary for effective management. Advanced sensor technologies, such as Coriolis and Magnetic Induction devices, offer higher accuracy but are cost-prohibitive for widespread adoption in economically constrained regions [3].
Previous research has demonstrated the promising potential of soft sensors for flow estimation based on fluid level [1, 3]. Specifically, Noori et al. [3] simulated drilling conditions using a flow circulation system with a Venturi structure, utilizing three level sensors for input data, while a Coriolis mass flow meter provided target data. A simple Multi-Layer Perceptron (MLP) model was then used to efficiently estimate non-Newtonian fluid flow.
Building on these datasets, our prior work [4] implemented a quantized MLP model on a low-power MCU (Cortex-M0 RP2040) and an embedded FPGA (Xilinx Spartan-7 XC7S15) using TensorFlow Lite and our customized ElasticAI.Creator toolchains [5], respectively. The FPGA-deployed models demonstrated up to 28.44$\times$ faster inference speed compared to MCU-deployed models, easily meeting real-time requirements. However, due to differences in applied quantization technologies, FPGA-deployed models exhibited lower precision. By applying integer-only quantization, we increased the model precision by up to 9.7% [6]. Additionally, implementing pipelined matrix multiplication on the FPGA reduced inference time by up to 9.39%, with an energy cost per inference only up to 5.71% higher.
Despite these advancements, there remains a need for datasets specifically for wastewater flow estimation. Additionally, the deployment of an integer-only quantized MLP model is not automated, making it challenging for researchers without FPGA expertise. Moreover, the previous hardware platform was overly powerful for this soft sensor application, leading to inefficient resource utilization and significant power consumption, resulting in energy inefficiency. This research explores end-to-end optimization strategies from sensor integration to model deployment, as shown in Figure $1$Some images are sourced from the internet, with all rights reserved by the original authors.. The goal is to develop a scalable, efficient system capable of real-time, accurate wastewater flow measurement, thereby enhancing in-field monitoring capabilities in managing urban overflows.
@@ II-C Optimized Hardware System Design
In this study, we also propose an optimized hardware design targeting energy-efficient soft sensor applications. The modification involves replacing the XC7S15 FPGA with a smaller FPGA, the ICE40UP5K. This decision is based on our previous work [6], which indicated low resource utilization of our MLP accelerator on the XC7S15 FPGA (6.47% LUTs, 7.5% BRAMs, and 10% DSPs). This low utilization suggests that the design can be shifted to a smaller FPGA without performance loss. Additionally, the XC7S15 FPGA has significant static power consumption (approximately 30 mW), whereas the ICE40UP5K’s static power consumption is at the micro-watt level. Thus, this replacement could markedly improve the system’s energy efficiency. To address the challenges posed by resource constraints, we plan to implement advanced hardware optimizations in the accelerator design. These optimizations will ensure that the design remains efficient and effective despite the reduced resources available on the ICE40UP5K.

### v2

@@ I Introduction and Related Work
Combined sewer systems (CSS) are widely used worldwide but face increasing challenges from extreme weather exacerbated by pollution and climate change. These conditions often lead to urban overflow events, significantly impacting public health and environmental sustainability [ahm2016estimation]. Real-time control systems offer a promising alternative by preventing CSS overflow through dynamic flow monitoring. However, these systems are hindered by high operational costs and challenges in achieving continuous, accurate monitoring and real-time data processing [sun2020integrated]. Current flow monitoring techniques mainly rely on hydrological formulas, which, while cost-effective, lack the precision necessary for effective management. Advanced sensor technologies, such as Coriolis and Magnetic Induction devices, offer higher accuracy but are cost-prohibitive for widespread adoption in economically constrained regions [noori2020non].
Previous research has demonstrated the promising potential of soft sensors for flow estimation based on fluid level [ahm2016estimation, noori2020non]. Specifically, Noori et al. [noori2020non] simulated drilling conditions using a flow circulation system with a Venturi structure, utilizing three level sensors for input data, while a Coriolis mass flow meter provided target data. A simple Multi-Layer Perceptron (MLP) model was then used to efficiently estimate non-Newtonian fluid flow.
[…]
Despite these advancements, there remains a need for datasets specifically for wastewater flow estimation. Additionally, the deployment of an integer-only quantized MLP model is not automated, making it challenging for researchers without FPGA expertise. Moreover, the previous hardware platform was overly powerful for this soft sensor application, leading to inefficient resource utilization and significant power consumption, resulting in energy inefficiency. This research explores end-to-end optimization strategies from sensor integration to model deployment, as shown in Figure 1111Some images are sourced from the internet, with all rights reserved by the original authors.. The goal is to develop a scalable, efficient system capable of real-time, accurate wastewater flow measurement, thereby enhancing in-field monitoring capabilities in managing urban overflows.
@@ II-C Optimized Hardware System Design
In this study, we also propose an optimized hardware design targeting energy-efficient soft sensor applications. The modification involves replacing the XC7S15 FPGA with a smaller FPGA, the ICE40UP5K. This decision is based on our previous work [ling2024flow], which indicated low resource utilization of our MLP accelerator on the XC7S15 FPGA (6.47% LUTs, 7.5% BRAMs, and 10% DSPs). This low utilization suggests that the design can be shifted to a smaller FPGA without performance loss. Additionally, the XC7S15 FPGA has significant static power consumption (approximately 30 mW), whereas the ICE40UP5K’s static power consumption is at the micro-watt level. Thus, this replacement could markedly improve the system’s energy efficiency. To address the challenges posed by resource constraints, we plan to implement advanced hardware optimizations in the accelerator design. These optimizations will ensure that the design remains efficient and effective despite the reduced resources available on the ICE40UP5K.